# A Survey on LLM-Assisted Clinical Trial Recruitment

### v2

@@ 7 Evaluation
Standard IR metrics, such as, precision, recall, F1, precision at rank k, normalized discounted cumulative gains (NDCG@k) and mean reciprocal rank (MRR), are the most popular evaluation metrics. The 2018 N2C2 cohort selection and TREC CT benchmarks have high performance metrics (Table 1), which are further improved by newer methods. Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data Kim and Quintana (2022); Kantor and Morzy (2024b).
Besides accuracy metrics, Rybinski et al. (2024) report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10. Jin et al. (2024) report a 42% reduction in screening time when using LLM-generated predictions and explanations. The measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent.
@@ 7.1 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we often come across inconsistent and semantically overlapping categories of errors in LLM generated explanations. For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by Jin et al. (2024); Hamer et al. (2023); Nievas et al. (2024), even though the latter often leads to the former (full list in Appendix B). In the proposal by Liévin et al. (2024), reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. Here, we identify where errors occur and break them down further by an identifiable source of the error (see Figure 2).
@@
These are errors in information extraction from patient data or trial criteria falls. This includes missed or incorrect NER, measurements (numeric or unit), temporal scopes and negations. LLMs are good at recognizing entities and measurements, but still struggle with negations Nievas et al. (2024).
@@
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities Wei et al. (2022). Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in Nievas et al. (2024)). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to evaluate. This shortcut, similar to cognitive biases in humans, hints towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
@@ 8 Discussion
Despite attempts to structure clinical trials Chen et al. (2022), finding similar trials and analyzing systematic failure cases is notoriously difficult Rybinski et al. (2020); White et al. (2023). Criterion-level annotations require significant manual annotations, thereby limiting the size of the corpus. For instance, Chia Kury et al. (2020) and LCT Dobbins et al. (2022), are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE Kang et al. (2017) and Criteria2Query Yuan et al. (2019), require expensive manual labels (230 disease-specific clinical trials in this case). Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy Kim and Quintana (2022). Systems thus evaluated on private retrospective data Wong et al. (2023); Yuan et al. (2023); Unlu et al. (2024) cannot be transparently compared.Kantor and Morzy (2024b) stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies. Public annotations, such as the TREC CT tracks have an average of 700 trial annotations per patient for less than 100 patients, while the N2C2 has only a few thousands of criterion-level annotations. Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by Kim and Quintana (2022) to generate large-scale data using automated methods.
@@
The benchmarks in Table 1 focus on model accuracy, corroborating the result from Bedi et al. (2025), which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured. Omar et al. (2024) reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined. Further results from Nemati et al. (2025), who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
@@
Equation 6.1 highlights the importance of aggregation and priority. Surprisingly, very few works explicitly formalize the matching task. This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions Hamer et al. (2023); Wornow et al. (2025); Unlu et al. (2024) and others. Formalization guides Yuan et al. (2023) to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, Jullien et al. (2024) and Saeidi (2025) to define re-ranking scores.
@@
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) Zhao et al. (2024b) and found that GPT4 tends to stereotype demographic presentations when generating diagnoses Zack et al. (2024). Alber et al. (2025) show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by Hamer et al. (2023), making the generalizability of LLMs across diseases unclear.
@@
While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations Unlu et al. (2024); Wong et al. (2023); Wornow et al. (2025), processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance Edemekong et al. (2024). As already discussed, creating large-scale realistic patient records while protecting their privacy is particularly challenging. There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative Das et al. (2023); Wang et al. (2024).
@@
According to Harrer (2023) the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7.1, we discussed how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization Guidance (2021).
@@ 9 Future Directions
As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention Gao et al. (2024); Qiu et al. (2024), we focus on four promising directions.
@@
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This problem has seen little activity since clustering using lexical features Hao et al. (2014). Newer search methods include constructing a clinical trial knowledge graph Chen et al. (2022), searching via patient EHRs Wu et al. (2018) and designing features for similarity matchingSun et al. (2022). Jullien et al. (2023) use textual entailment in LLMs to find trials that match short descriptions. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
@@
LLM agents have the potential of bridging the semantic gap between eligibility criteria and patient data by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new Luo et al. (2013), yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models Kang et al. (2017); Sun and Loparo (2019); Liu et al. (2021a); Dasgupta et al. (2020). Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria Clinical Trials Arena (2022). LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate Liu et al. (2021b).
@@
Liu et al. (2025c) propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction. Markey et al. (2025) showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance. Similar to Shi et al. (2024), who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result. Another important operation is learning to defer to experts Mozannar and Sontag (2020), which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
@@
Explaining black-box LLM predictions in human-understandable form is very challenging Zhao et al. (2024a). Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability Nauta et al. (2023); Bodria et al. (2023); Chen and Eickhoff (2024). Wong et al. (2023) could potentially explain eligibility prediction via logical component (mis)matches. Furthermore, the utility of these explanations are unclear, as they have been only evaluated manually. We hope that the error taxonomy discussed in Section 7.1 assists in systematic evaluation of explanations.
@@ Appendix A Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community, with standardized reporting guidelines Moher et al. (2009); Page et al. (2021). This guideline has a checklist of required items in the title, abstract, introduction, methods, results, discussion and other information that every study needs to fulfil. These studies are bound by strict inclusion and exclusion criteria applied to the scientific literature considered for any assessment. This rigor is necessary in evidence-based analysis of a specific research question. A survey is more flexible and provides a board coverage of the topic being discussed. For instance, the guidelines for writing surveys in our community as outlined by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
@@ Appendix B Reported Error Types
Errors reported in Jin et al. (2024) supplementary:
@@
Errors reported in Hamer et al. (2023):
@@
Errors reported in Nievas et al. (2024):
@@ Appendix C Trial-Patient Matching Systems
Direction Method Patient Source Trial Source Limitations Trial-centric Penberthy et al. (2010) Discrete-data filter and sub-word matching Availability:
@@
Yuan et al. (2019) (Criteria2Query)
@@
Criterion-level prediction Hamer et al. (2023)
@@
Trial-level prediction Wong et al. (2023)
@@
Yuan et al. (2023) Prompt LLMs to reformulate eligibility criteria Availability: Private
@@
Trial ranking Pradeep et al. (2022)
@@
Cosine similarity between patient and trial embeddings obtained using GPT by Richmond and Deshpande (2023) and using Sentence Transformer by Lahiri et al. (2023). Availability: Public (TREC CT 2023) Very low MAP of 0.02 Richmond and Deshpande (2023) and close to zero for Lahiri et al. (2023)
Kusa et al. (2023b)
Sentence query formulation using GPT-3.5. Query enrichment using Kusa et al. (2023a). Trial-level prediction of re-ranked pairs using GPT-3.5.
@@
Rybinski and Karimi (2023); Rybinski et al. (2024)
@@
Jin et al. (2024)
Hybrid trial filtering with BM25 and MedCPT Jin et al. (2023)
@@
Availability: Public (Koopman and Zuccon (2016), TREC CT 2021, 2022)
@@
Nievas et al. (2024)
Extends Jin et al. (2024) to open-source LLMs
(same as Jin et al. (2024))
@@
Saeidi et al. (2023); Saeidi (2025) Embed patients and trials to concept vector space. Use variations of Equation 6.1 to compute relevance scores. Availability: Public (TREC CT 2023) No direct connection between criterion-level LLM predictions and trial-level relevance score computations

### v3

@@ 7 Evaluation
Standard IR metrics, such as, precision, recall, F1, precision at rank k, normalized discounted cumulative gains (NDCG@k) and mean reciprocal rank (MRR), are the most popular evaluation metrics. The 2018 N2C2 cohort selection and TREC CT benchmarks have high performance metrics (Table 1), which are further improved by newer methods. Notably, researchers in the medical domain have stressed the lack of consistent benchmark data, primarily, due to the dependence on manual review which also restricts the size of the data kim2022review; kantor2024machine.
Besides accuracy metrics, rybinski2024learning report query latency, with GPT models having 15 times higher latency (47.8s) when compared to smaller pre-trained models for a 12% increase in NDCG@10. jin2024matching report a 42% reduction in screening time when using LLM-generated predictions and explanations. The measurement of human preference of LLM generations, bias in model predictions and the informativeness of LLM explanations ranges from limited to non-existent.
@@ 7.1 Taxonomy of LLM Errors
Since there is no one accepted taxonomy of errors, we often come across inconsistent and semantically overlapping categories of errors in LLM generated explanations. For e.g., incorrect reasoning and lack of knowledge are recognized as independent error types by jin2024matching; hamer2023improving; nievas2024distilling, even though the latter often leads to the former (full list in Appendix B). In the proposal by lievin2024can, reasoning errors are separate from reading comprehension, even though the instances of incorrect reading comprehension occur when the model ignores contextual information and “reasons” using its learned knowledge. Here, we identify where errors occur and break them down further by an identifiable source of the error (see Figure 2).
@@
These are errors in information extraction from patient data or trial criteria falls. This includes missed or incorrect NER, measurements (numeric or unit), temporal scopes and negations. LLMs are good at recognizing entities and measurements, but still struggle with negations nievas2024distilling.
@@
Generating explanations to arrive at final answers unlock LLM’s reasoning capacities wei2022chain. Even so, the explanation and the prediction can be inconsistent. LLMs may predict incorrectly despite correct explanation (reported as explanation-output mismatch in nievas2024distilling). The opposite situation, when a prediction is correct despite an incorrect explanation, is more difficult to evaluate. This shortcut, similar to cognitive biases in humans, hints towards a bias that the model picked up during training. Both these cases negatively impact transparency and accountability of the matching system.
@@ 8 Discussion
Despite attempts to structure clinical trials chen2022knowledge, finding similar trials and analyzing systematic failure cases is notoriously difficult rybinski2020clinical; white2023clinidigest. Criterion-level annotations require significant manual annotations, thereby limiting the size of the corpus. For instance, Chia kury2020chia and LCT dobbins2022leaf, are manually annotated corpora of 1000 trials each, while supervised annotators such as, EliIE kang2017eliie and Criteria2Query yuan2019criteria2query, require expensive manual labels (230 disease-specific clinical trials in this case). Eligibility labels on real patient data from real enrollment status, necessary to discount reviewer bias, are only available on a small-scale, if at all, due to the trade-off between scale and privacy kim2022review. Systems thus evaluated on private retrospective data wong2023scaling; yuan2023large; unlu2024retrieval cannot be transparently compared.kantor2024machine stress on standardized benchmarks as the dependence on manual evaluation hinders meta-analyses and comparison between different studies. Public annotations, such as the TREC CT tracks have an average of 700 trial annotations per patient for less than 100 patients, while the N2C2 has only a few thousands of criterion-level annotations. Jointly, the major public corpora, e.g., ClinicalTrials.gov and MIMIC datasets, present an opportunity to build on the proposal by kim2022review to generate large-scale data using automated methods.
@@
The benchmarks in Table 1 focus on model accuracy, corroborating the result from bedi2025testing, which reports that more than 95% of studies use accuracy as the primary dimension of evaluation, while fairness, bias and uncertainty are less frequently measured. omar2024large reviewed 27 clinical trials evaluating LLMs in healthcare also found the accuracy and reliability standards for LLM use to be undefined. Further results from nemati2025benchmarking, who benchmarked the annotation ability of LLMs across 9 performance metrics, show that while LLMs consistently score high on precision, recall and F1 (lowest being 0.8), their scores highly vary on semantic similarity, factual consistency, relevance, fluency, consistency and coherence (ranging from 0.1 to 0.9) highlighting the need for multiple dimensions of evaluation.
@@
Equation 6.1 highlights the importance of aggregation and priority. Surprisingly, very few works explicitly formalize the matching task. This lack of formalization coincides with an absence of aggregation strategies for trial-level predictions hamer2023improving; wornow2025zero; unlu2024retrieval and others. Formalization guides yuan2023large to design a loss function that accounts for the contrastive requirements of inclusion and exclusion criteria, jullien2024controlled and saeidi2025streamlining to define re-ranking scores.
@@
Recent research measured distinct bias in disease diagnosis across gender, age and disease in popular LLM models (GPT4, ChatGPT and Qwen) zhao2024can and found that GPT4 tends to stereotype demographic presentations when generating diagnoses zack2024assessing. alber2025medical show that LLMs are prone to making medical misjudgments by replacing just 0.001% of the training data with medical misinformation. All the LLM-based systems, except for the trial ranking models, are evaluated on a few disease-specific trials (see Table LABEL:tab:systems_overview_llm in Appendix C), the largest being 146 trials, covering 10 cancer types, evaluated by hamer2023improving, making the generalizability of LLMs across diseases unclear.
@@
While we found several methods that deployed GPT models on Azure AI to comply with privacy regulations unlu2024retrieval; wong2023scaling; wornow2025zero, processing patient data with LLMs still raises serious ethical challenges due to the lack of HIPAA compliance edemekong2024health. As already discussed, creating large-scale realistic patient records while protecting their privacy is particularly challenging. There is interest in generating synthetic data, called digital twins, with limited access to real patient data as a viable privacy protected alternative das2023twin; wang2024twin.
@@
According to harrer2023attention the core limitations affecting LLM adoption are unfiltered pre-training, which does not differentiate between facts, opinions, or misinformation; lack of self-assessment, where a model generates invalid but syntactically and semantically coherent sentences; non-determinism: where surface-form prompt variations lead to drastic changes in the output and repeatability is not guaranteed under consistent input conditions; and, knowledge recall: where updating outdated data or injecting new information requires expensive retraining since the mechanisms of memory in LLMs are not well understood. In Section 7.1, we discussed how knowledge recall and lack of self-assessment surface through reasoning errors. These limitations pose a direct challenge to the transparency and accountability principles of AI for health laid down by the World Health Organization guidance2021ethics.
@@ 9 Future Directions
As interest in LLMs orchestrating an end-to-end pipeline and incorporating human interactions is gaining more attention gao2024empowering; qiu2024llm, we focus on four promising directions.
@@
Past trials that share a target population are important for designing new trials, recruiting patients, systematic reviews and meta-analyses. This problem has seen little activity since clustering using lexical features hao2014clustering. Newer search methods include constructing a clinical trial knowledge graph chen2022knowledge, searching via patient EHRs wu2018semehr and designing features for similarity matchingsun2022interactivess. jullien2023nli4ct use textual entailment in LLMs to find trials that match short descriptions. However, these models falter on inferences that require numerical reasoning and could not surpass a BM25 baseline for ranking evidence.
@@
LLM agents have the potential of bridging the semantic gap between eligibility criteria and patient data by suggesting data models underlying patient data for structuring eligibility criteria early on. This collaborative idea is not new luo2013human, yet, designing eligibility criteria is a big challenge and has been handled post-hoc by optionally considering patient data models kang2017eliie; sun2019knowledge; liu2021knowledge; dasgupta2020extracting. Small patient pools, which ultimately affect the successful completion of a trial, are often the result of restrictive criteria clinicaltrialarena2022. LLMs can improve trial design by identifying restrictive criteria for trial investigators to relax them and create larger patient pools, especially for trials tackling diseases with a high mortality rate liu2021evaluating.
@@
liu2025autoct propose an iterative feature discovery model using LLM agents for interpretable trial outcome prediction. markey2025rags showed promising results on content relevance and suitability of trial protocols generated using LLMs, with room for improvement in logical reasoning and provenance. Similar to shi2024enhancing, who propose collaboration of agents for knowledge-augmentation and reasoning, LLM agents can identify and distribute tasks and aggregate them to a final result. Another important operation is learning to defer to experts mozannar2020consistent, which can separate operable criteria, such as those that require tabular operations (e.g., via structured queries) or reasoning (temporal, numerical, negation), from criteria that require expert feedback.
@@
Explaining black-box LLM predictions in human-understandable form is very challenging zhao2024explainability. Explainability in clinical trial matches is limited to chain-of-thought generations, which is only one of the many facets of explainability nauta2023anecdotal; bodria2023benchmarking; chen2024evaluating. wong2023scaling could potentially explain eligibility prediction via logical component (mis)matches. Furthermore, the utility of these explanations are unclear, as they have been only evaluated manually. We hope that the error taxonomy discussed in Section 7.1 assists in systematic evaluation of explanations.
@@ Appendix A Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community, with standardized reporting guidelines moher2009preferred; page2021prisma. This guideline has a checklist of required items in the title, abstract, introduction, methods, results, discussion and other information that every study needs to fulfil. These studies are bound by strict inclusion and exclusion criteria applied to the scientific literature considered for any assessment. This rigor is necessary in evidence-based analysis of a specific research question. A survey is more flexible and provides a board coverage of the topic being discussed. For instance, the guidelines for writing surveys in our community as outlined by ACMCSUR, TACL and ARR aim to draw perspectives on an evolving topic of interest.
@@ Appendix B Reported Error Types
Errors reported in jin2024matching supplementary:
@@
Errors reported in hamer2023improving:
@@
Errors reported in nievas2024distilling:
@@ Appendix C Trial-Patient Matching Systems
Direction Method Patient Source Trial Source Limitations Trial-centric penberthy2010automated Discrete-data filter and sub-word matching Availability:
@@
yuan2019criteria2query (Criteria2Query)
@@
Criterion-level prediction hamer2023improving
@@
Trial-level prediction wong2023scaling
@@
yuan2023large Prompt LLMs to reformulate eligibility criteria Availability: Private
@@
Trial ranking pradeep2022neural
@@
Cosine similarity between patient and trial embeddings obtained using GPT by richmond2023leveraging and using Sentence Transformer by lahiri2023tmu. Availability: Public (TREC CT 2023) Very low MAP of 0.02 richmond2023leveraging and close to zero for lahiri2023tmu
kusa2023dossier
Sentence query formulation using GPT-3.5. Query enrichment using kusa2023effective. Trial-level prediction of re-ranked pairs using GPT-3.5.
@@
rybinski2023matching; rybinski2024learning
@@
jin2024matching
Hybrid trial filtering with BM25 and MedCPT jin2023medcpt
@@
Availability: Public (koopman2016test, TREC CT 2021, 2022)
@@
nievas2024distilling
Extends jin2024matching to open-source LLMs
(same as jin2024matching)
@@
saeidi2023malnis; saeidi2025streamlining Embed patients and trials to concept vector space. Use variations of Equation 6.1 to compute relevance scores. Availability: Public (TREC CT 2023) No direct connection between criterion-level LLM predictions and trial-level relevance score computations